OpenStudy (anonymous):

using substitution: 4v-2u=9 2v-4u=-3

OpenStudy (anonymous):

u = 5/2 v = 7/2

OpenStudy (anonymous):

4v-2u=9 v = (9+2u)/4

OpenStudy (anonymous):

2v-4u=-3 2((9+2u)/4 )-4u = -3

OpenStudy (anonymous):

can u solve now?

OpenStudy (anonymous):

(18+4u)/4 = 4u -3 4u + 18 = 16u - 12

OpenStudy (anonymous):

12u = 30 u = 30/12=5/2 v = (9+2u)/4 = (9+10/2)/4 (28/2)/4 28/2 X 1/4 7/2

OpenStudy (anonymous):

complete?